# IFT-UAM/CSIC-24-127

### v1

@@ 1 Introduction
Despite this recent progress, there is still no firm understanding of how the near-singularity Kasner exponents and their potentially chaotic dynamics arise from a boundary perspective. In order to isolate the relevant bulk dynamics, and to be able to fully explore its possibilities, we propose a new approach to the near-singularity dynamics of general relativity coupled to matter, focusing on the ultra-local structure that arises naturally at late interior times. In this limit, the light cone collapses to a line, which corresponds to a ‘small speed of light’ contraction of the Lorentz algebra that leads to the Carroll algebra [39, 40]. Geometrically, the Carroll limit can be described in terms of fully covariant curved spacetime geometry using a toolkit similar to the Newton–Cartan geometry associated to non-relativistic limits, as recently reviewed in [41].
@@
In this paper, however, we will use the ultra-local Carroll limit to describe bulk physics, focusing specifically on the near-singularity BKL dynamics of black holes. Our main aim will be to demonstrate that a theory of Carroll gravity, consisting of dynamical Carroll geometry coupled to matter, can describe the aforementioned mixmaster behavior. Moreover, our Carroll theory of gravity is directly obtained from an ultra-local limit of the bottom-up AdS/CFT model that was recently introduced in [23] to obtain mixmaster dynamics near the singularity of an asymptotically AdS black hole.
@@
We want to argue that this new approach to near-singularity dynamics is promising for several reasons. First, since the ultra-local limit is implemented off shell on the level of the geometry, the Carroll theory of gravity that we consider [49, 50, 51, 52] is inherently much simpler than full general relativity (GR). In particular, as emphasized in [52, 53], its evolution equations always consist of ordinary differential equations, instead of the hyperbolic partial differential equations of GR. As such, our approach will allow us to construct a much larger class of models where the emergence of BKL-type behavior can be observed explicitly, either analytically or numerically, than what would be tractable in full GR. Additionally, the Carroll theory of gravity that we use can be seen as the leading order in the ultra-local expansion of GR that was recently constructed in [52], building on a related construction of the off-shell non-relativistic expansion in [54, 55, 56, 57]. While we only focus on the leading-order action in the present paper, our approach is therefore naturally equipped to explicitly studying subleading corrections to near-singularity BKL dynamics in a tractable way.
@@
This paper is organized as follows. We start out by reviewing salient features of BKL dynamics in Einstein gravity. For this, we first introduce a useful parametrization of Kasner-type geometries with time-dependent scaling exponents in Section 2.1, and we show how it gives rise to the interpretation of such geometries as trajectories in a minisuperspace parametrized by the scaling exponents. In Section 2.2, we then introduce the mixmaster model, and we show how it leads to nontrivial trajectories, which can furthermore be mapped to billiard dynamics on a hyperbolic triangle. Section 3 then introduces the necessary Carroll tools, including a brief overview of curved Carroll geometry, the off-shell ultra-local expansion of general relativity introduced in [52] and the coupling to Carroll limits of matter actions. We then show in Section 4.1 that leading-order vacuum Carroll gravity admits Kasner solutions [^1] The fact that Carroll gravity has Kasner-type solutions with fixed scaling exponents as in (1.1) was observed in [58, 59] and was also pointed out by M. Henneaux during the 2022 Vienna Carroll Workshop. However, BKL or mixmaster dynamics has not yet been reproduced from Carroll gravity. and similarly leads to trajectories in an external Minkowski space. Motivated by the recently-proposed AdS/CFT model from [23], we then show in Section 4.2 that the Carroll limit of gravity coupled to three abelian gauge fields reproduces the full dynamics of the mixmaster model. Finally, we summarize our findings and set out several future directions in Section 5.
@@ 3 Leading-order Carroll gravity
As we will see, the tensor $h^{\mu\nu}$ can be interpreted as the inverse of a (degenerate) spatial metric $h_{\mu\nu}$ which, together with a ‘time’ vector field $v^{\mu}$ determines the leading-order Carroll geometry. [^3] Since the spatial metric $h_{\mu\nu}$ is degenerate, the inverse $h^{\mu\nu}$ is not uniquely determined, which is reflected by its transformation (3.7) under local Carroll boosts. Since the leading-order Carroll limit of the Maxwell action only contains the electric field, it is often called the ‘electric’ limit of electromagnetism [60, 51, 61, 59]. By analogy, the leading-order Carroll limit (3.1) obtained from the Einstein–Hilbert action is also sometimes known as the ‘electric Carroll limit’ of general relativity. [^4] Using different parametrizations, ‘magnetic’ Carroll limits can also be obtained, but we will not need them here. Whereas electric actions typically only contain time derivatives of the fields (corresponding to the extrinsic curvature in gravity), magnetic Carroll actions typically contain only spatial derivatives (or spatial curvature in gravity). To retain Carroll invariance, the latter also come with constraints that strongly restrict their dynamics in time, so they are unsuitable for describing BKL dynamics.
@@ 3.3 Carroll gravity coupled to electric limit of Maxwell
Now let us consider the specific example of the Carroll limit of gravity coupled to electromagnetism. Following a similar procedure as what we outlined above for the Einstein–Hilbert action, the leading-order term in the ultra-local $c\to 0$ limit of the Maxwell action leads to the ‘electric’ Carroll limit of electromagnetism [60, 59]
@@ 5 Summary and outlook
In this work, we have put forward the ultra-local Carroll expansion of general relativity as a new and useful tool for studying Belinski–Khalatnikov–Lifshitz (BKL) dynamics near spacelike singularities. As a first example, we have explicitly shown that mixmaster dynamics can be obtained from the leading-order Carroll limit of gravity coupled to three abelian gauge fields, following the bottom-up AdS/CFT model which was introduced recently in [23]. This further suggests that Carroll expansions may be a useful tool in constructing new tractable models of the dynamics beyond the horizon in holography.
@@
Next, the Carroll gravity models we employed above are in fact only the leading order theories in a systematic ultra-local expansion of general relativity [52]. What is more, the intrinsic curvature of spatial slices does not enter into this expansion until next-to-leading order. As we illustrated in Section 2.2, spatial curvature plays a key role in BKL dynamics by sourcing potential walls such as in the original $SO(3)$ or Bianchi IX mixmaster model [3]. In the above, following [23], we were able to reproduce dynamics equivalent to the mixmaster model using only potential walls sourced by matter. However, to capture the full richness of BKL dynamics, our current Carroll models should be extended to also be sensitive to spatial curvature and its resulting potential walls.
@@ Acknowledgements
We are thankful to Ján Pulmann and Dmitriy Zhigunov for useful discussions, and to Jelle Hartong for useful discussions and helpful comments on an early version of this paper. GO would like to thank the organizers and participants of the ‘Eurostrings 2024 meets Fundamental Physics UK’ conference (Southampton) and the 2024 ‘Carrollian Physics and Holography’ workshop (Erwin Schrödinger Institute, Vienna), where part of this work was presented.

### v2

@@ 1 Introduction
Despite this recent progress, there is still no firm understanding of how the near-singularity Kasner exponents and their potentially chaotic dynamics arise from a boundary perspective. In order to isolate the relevant bulk dynamics, and to be able to fully explore its possibilities, building on earlier observations in [39, 4], we develop a novel approach to the near-singularity dynamics of general relativity coupled to matter, focusing on the ultra-local structure that arises naturally at late interior times. In this limit, the light cone collapses to a line, which corresponds to a ‘small speed of light’ contraction of the Lorentz algebra that leads to the Carroll algebra [40, 41]. Geometrically, the Carroll limit can be described in terms of fully covariant curved spacetime geometry using a toolkit similar to the Newton–Cartan geometry associated to non-relativistic limits, as recently reviewed in [42].
@@
In this paper, however, we will use the ultra-local Carroll limit to describe bulk physics, focusing specifically on the near-singularity BKL dynamics of black holes. Our main aim will be to demonstrate explicitly that a theory of Carroll gravity, consisting of dynamical Carroll geometry coupled to matter, can describe the aforementioned mixmaster behavior. Moreover, this Carroll theory of gravity is directly obtained from an ultra-local limit of the bottom-up AdS/CFT model that was recently introduced in [23] to obtain mixmaster dynamics near the singularity of an asymptotically AdS black hole.
@@
We want to argue that this approach to near-singularity dynamics is promising for several reasons. First, since the ultra-local limit is implemented off shell on the level of the geometry, the Carroll theory of gravity that we consider [50, 51, 52, 53] is inherently much simpler than full general relativity (GR). In particular, as emphasized in [50, 39, 54, 53], its evolution equations always consist of ordinary differential equations, instead of the hyperbolic partial differential equations of GR. As such, our approach will allow us to construct a much larger class of models where the emergence of BKL-type behavior can be observed explicitly, either analytically or numerically, than what would be tractable in full GR. Additionally, the Carroll theory of gravity that we use can be seen as the leading order in the ultra-local expansion of GR that was recently constructed in [53], building on a related construction of the off-shell non-relativistic expansion in [55, 56, 57, 58]. While we only focus on the leading-order action in the present paper, our approach is therefore naturally equipped to explicitly studying subleading corrections to near-singularity BKL dynamics in a tractable way.
@@
This paper is organized as follows. We start out by reviewing salient features of BKL dynamics in Einstein gravity. For this, we first introduce a useful parametrization of Kasner-type geometries with time-dependent scaling exponents in Section 2.1, and we show how it gives rise to the interpretation of such geometries as trajectories in a minisuperspace parametrized by the scaling exponents. In Section 2.2, we then introduce the mixmaster model, and we show how it leads to nontrivial trajectories, which can furthermore be mapped to billiard dynamics on a hyperbolic triangle. Section 3 then introduces the necessary Carroll tools, including a brief overview of curved Carroll geometry, the off-shell ultra-local expansion of general relativity introduced in [53] and the coupling to Carroll limits of matter actions. We then show in Section 4.1 that leading-order vacuum Carroll gravity admits Kasner solutions [^1] The fact that Carroll gravity has Kasner-type solutions with fixed scaling exponents as in (1.1) was observed in [59, 60] and was also pointed out by M. Henneaux during the 2022 Vienna Carroll Workshop. However, these references did not reproduce BKL or mixmaster dynamics from Carroll gravity. Our aim in the present paper is to demonstrate this connection explicitly, following earlier observations in [39, 4], and similarly leads to trajectories in an external Minkowski space. Motivated by the recently-proposed AdS/CFT model from [23], we then demonstrate explicitly in Section 4.2 that the Carroll limit of gravity coupled to three abelian gauge fields reproduces the full dynamics of the mixmaster model. Finally, we summarize our findings and set out several future directions in Section 5.
@@ 3 Leading-order Carroll gravity
As we will see, the tensor $h^{\mu\nu}$ can be interpreted as the inverse of a (degenerate) spatial metric $h_{\mu\nu}$ which, together with a ‘time’ vector field $v^{\mu}$ determines the leading-order Carroll geometry. [^3] Since the spatial metric $h_{\mu\nu}$ is degenerate, the inverse $h^{\mu\nu}$ is not uniquely determined, which is reflected by its transformation (3.7) under local Carroll boosts. Since the leading-order Carroll limit of the Maxwell action only contains the electric field, it is often called the ‘electric’ limit of electromagnetism [50, 61, 52, 62, 60]. By analogy, the leading-order Carroll limit (3.1) obtained from the Einstein–Hilbert action is also sometimes known as the ‘electric Carroll limit’ of general relativity. [^4] Using different parametrizations, ‘magnetic’ Carroll limits can also be obtained, but we will not need them here. Whereas electric actions typically only contain time derivatives of the fields (corresponding to the extrinsic curvature in gravity), magnetic Carroll actions typically contain only spatial derivatives (or spatial curvature in gravity). To retain Carroll invariance, the latter also come with constraints that strongly restrict their dynamics in time, so they are unsuitable for describing BKL dynamics.
@@ 3.3 Carroll gravity coupled to electric limit of Maxwell
Now let us consider the specific example of the Carroll limit of gravity coupled to electromagnetism. Following a similar procedure as what we outlined above for the Einstein–Hilbert action, the leading-order term in the ultra-local $c\to 0$ limit of the Maxwell action leads to the ‘electric’ Carroll limit of electromagnetism [50, 61, 60]
@@ 5 Summary and outlook
In this work, we have put forward the ultra-local Carroll expansion of general relativity as a novel and useful tool for studying Belinski–Khalatnikov–Lifshitz (BKL) dynamics near spacelike singularities. As a first example, following earlier observations in [39, 4], we have explicitly shown that mixmaster dynamics can be obtained from the leading-order Carroll limit of gravity coupled to three abelian gauge fields, following the bottom-up AdS/CFT model which was introduced recently in [23]. This further suggests that Carroll expansions may be a useful tool in constructing new tractable models of the dynamics beyond the horizon in holography.
@@
Next, the Carroll gravity models we employed above are in fact only the leading order theories in a systematic ultra-local expansion of general relativity [53]. What is more, the intrinsic curvature of spatial slices does not enter into this expansion until next-to-leading order. As we illustrated in Section 2.2, spatial curvature plays a key role in BKL dynamics by sourcing potential walls such as in the original $SO(3)$ or Bianchi IX mixmaster model [3]. In the above, building on earlier observations in [39, 4], we were able to reproduce dynamics equivalent to the mixmaster model in a Carroll limit of the model proposed in [23], using only potential walls sourced by matter. However, to capture the full richness of BKL dynamics, our current Carroll models should be extended to also be sensitive to spatial curvature and its resulting potential walls.
@@ Acknowledgements
We are thankful to Ján Pulmann and Dmitriy Zhigunov for useful discussions, and to Jelle Hartong and Marc Henneaux for useful discussions and helpful comments on an earlier version of this paper. GO would like to thank the organizers and participants of the ‘Eurostrings 2024 meets Fundamental Physics UK’ conference (Southampton) and the 2024 ‘Carrollian Physics and Holography’ workshop (Erwin Schrödinger Institute, Vienna), where part of this work was presented.